OpenStudy (anonymous):

A drugstore sells a drug costing 85 for 112 and a drug costing 175 for 238. If the markup policy of the drugstore is assumed to b e linear, write an equation the expresses retail price R in terms of cost C (wholesale price).

OpenStudy (nurali):

R = mC +d 112 = 85C + d (1) 238 = 175C +d (2) (2) -(1) gives 126 = 90C C = 126/90 = 1 .4 Therefore R = 1.4C + d Sub. in (85,112) 112 = 1.4(85)+ d d = 112 -1.4(85) = 112 - 119 = -7 So equation is R = 1.4C -7

OpenStudy (anonymous):

Thanks, Nurali! Appreciated.